plz help!!!


A merry-go-round with 5 m radius is rotating around its vertical axis with a rate of 0.30 rev/s. A 90 kg man jump on to the merry-go-round and he lands and remains at a distance of 3 m from the axis of rotation. If the moment of inertia of the merry-go-round is 10,000

kgm2, find the new rate of rotation after the person has landed.


1
Expert's answer
2021-09-22T07:05:29-0400

Given:

"r=5\:\rm m"

"n=30\:\rm rev/s"

"I=10000\:\rm kg\cdot m^2"

"m=90\;\rm kg"

"d=3\:\rm m"


The law of conservation of angular momentum says

"In=I'n'"

"I'=I+md^2"

Hence

"n'=\frac{I}{I+md^2}n=\frac{10000}{10000+90*3^2}*30=28\:\rm rev/s"